OpenStudy (anonymous):

An Auditorium has 20 rows of seats. The front row has 45 seats. Each row after that has 4 more seats than the previous row. FInd the total Number of seats in the auditorium. !!!!Please need help idk how to do this plz!!!

OpenStudy (anonymous):

It's a arimethic series question.

OpenStudy (anonymous):

\[S _{n}=n/2(t _{1}+t _{n})\]

OpenStudy (anonymous):

\[t _{n}=t _{1}+(n-1)d\]

OpenStudy (anonymous):

1st row = 45 seats. Common difference is 4.

OpenStudy (anonymous):

The 20th term, Tsubscript(20) = 45 + (20-1)4 Tsubscript(20) = 121. Back to series formula Tsubscript(20) = 20/2(45+121) = 1660 total seats. @torken8

OpenStudy (anonymous):

@torken8 Am I right?

OpenStudy (anonymous):

yea thx much man